# Unveiling m-Sharpness Through the Structure of Stochastic Gradient Noise

### v2

@@ 5.1 Training from scratch
We evaluate three optimization methods SGD, SAM [^2] In this section, consistent with common practice, we use “SAM” to refer to the mini-batch SAM., and RW-SAM on CIFAR-10 and CIFAR-100 (Krizhevsky et al.,, 2009), training three models from scratch: ResNet-18, ResNet-50 (He et al.,, 2016), and WideResNet-28-10 (Zagoruyko and Komodakis,, 2016). We use a batch size of 128 and a cosine learning rate schedule with an initial learning rate of 0.1. SAM and RW-SAM are trained for 200 epochs, while SGD is trained for 400 epochs. We apply a momentum of 0.9 and a weight decay of $5e^{-4}$, along with standard data augmentation techniques, including horizontal flipping, padding by four pixels, and random cropping.
@@
For large-scale experiments, we train a ResNet-50 on ImageNet-1K (Deng et al.,, 2009) for 90 epochs with an initial learning rate of 0.05. For both SAM and RW-SAM, we use the same $\rho=0.05$. For the additional hyperparameter $\lambda$ in RW-SAM, we set $\lambda=0.25$. All other hyperparameters remain the same as those used on CIFAR-10/100.
@@
Analysis of Computational Overhead. RW-SAM requires an additional forward pass to estimate per-sample gradient norms, leading to approximately 1/6 more training overhead compared to vanilla SAM, as a forward pass typically takes about half the time of a backward pass (Kaplan,, 2022). For a report of the additional wall‐clock time overhead, please see Table 8 in Appendix J. However, RW-SAM matches the performance of m-SAM at $m=64$ without incurring its nearly two-fold training overhead (See Table 7 in Appendix J). This highlights the efficiency of RW-SAM in balancing computational cost and performance.
@@ 5.2 Fine-tuning
We fine-tune a ViT-B/16 model (Dosovitskiy et al.,, 2020), pre-trained on ImageNet-1K, on CIFAR-10 and CIFAR-100. We train for 20 epochs with an initial learning rate of 0.01. Other hyperparameters are the same as those in training from scratch. The results are summarized in Table 2(b). We also fine-tune a pretrained DistilBERT model (Sanh et al.,, 2019) on the GLUE benchmark (Wang et al.,, 2018). We use AdamW (Loshchilov and Hutter,, 2019) as the base optimizer. The detailed hyperparameter settings are provided in Table 10 of Appendix J, and the results are presented in Table 3. We observe that RW-SAM consistently outperforms SAM in both fine-tuning experiments.
@@ 5.3 Robustness to label noise
Foret et al., (2021) have shown that SAM exhibits robustness to label noise. Motivated by this, we evaluate RW-SAM’s performance on CIFAR-10 with labels randomly flipped at specified noise ratios. We train a ResNet-18 and report clean test accuracies in Table 4. As the noise ratio increases, RW-SAM maintains remarkably strong performance. In particular, at an 80 % noise ratio, RW-SAM achieves a 16 % absolute accuracy improvement over SAM.
@@ Acknowledgement
Following the SDE framework by Mil’shtein, (1986); Li et al., (2017); Compagnoni et al., (2023); Luo et al., (2025), we have the following definition:
@@
(Adaption of Theorem 3 in Li et al., (2019)) Let $T>0$, $\eta\in(0,\min\{1,T\})$, $N=\lfloor T/\eta\rfloor$. Let $\alpha\geq 1$ be an integer. Suppose further that the following conditions hold:
@@
By substituting Assumption (i) in the penultimate step of the proof in Theorem 3 in Li et al., (2019) with Assumption (i) of our Theorem A.2, the same argument goes through and yields the desired conclusion. Hence, we omit the proof.
Next, we state the key lemma for the two-parameter weak approximation. By employing a Dynkin expansion (see, e.g. Evans, (2012)) instead of a full Itô–Taylor expansion, it avoids the proliferation of terms and provides a streamlined approach to controlling the remainder error in the two-parameter setting.
@@ B.1 Mini-batch USAM
Theorem B.2 (mini-batch USAM SDE, adapted from Theorem 3.2 of Compagnoni et al., (2023)).
@@
Theorem B.2 follows by replacing the single‐parameter Lemmas A.1, A.2 and A.5 in Compagnoni et al., (2023) with our two‐parameter versions—Theorem A.2, Lemma A.4 and Lemma A.5—imposing the extra global Lipschitz conditions to guarantee existence and uniqueness of the strong solution, and using our Assumption 3.1 to expand the drift term. We therefore omit the routine algebraic details. ∎
@@
By Taylor’s theorem with integral remainder (Folland,, 2005) we have, for each $\gamma$,
@@
The Lipschitz conditions are to ensure that the SDE has a unique strong solution with uniformly bounded moments. It is possible to appropriately relax them if we allow weak solutions (Mil’shtein,, 1986).
@@
By Taylor’s theorem with integral remainder (Folland,, 2005) we have, for each $\gamma$,
@@
The Lipschitz conditions are to ensure that the SDE has a unique strong solution with uniformly bounded moments. It is possible to appropriately relax them if we allow weak solutions (Mil’shtein,, 1986).
@@ C.1 Mini-batch SAM
Theorem C.2 (mini-batch SAM SDE, adapted from Theorem 3.5 of Compagnoni et al., (2023)).
@@
    $\nabla f^{\mathrm{SAM}}\;=\;\nabla\Bigl(f+\rho\mathbb{E}\|\nabla f_{\gamma}\|)\quad\text{and}\quad\sqrt{\Sigma^{\mathrm{SAM}}}$
@@
Theorem C.2 follows by replacing the single‐parameter Lemmas A.1, A.2 and A.14 in Compagnoni et al., (2023) with our two‐parameter versions—Theorem A.2, Lemma A.4 and Lemma A.5—imposing the extra global Lipschitz conditions to guarantee existence and uniqueness of the strong solution. We therefore omit the routine algebraic details. ∎
@@
By Taylor’s theorem with integral remainder (Folland,, 2005), for each $\gamma$ and writing $u(x)=\nabla f(x)/\|\nabla f(x)\|$, we have
@@
The Lipschitz conditions are to ensure that the SDE has a unique strong solution with uniformly bounded moments. It is possible to appropriately relax them if we allow weak solutions (Mil’shtein,, 1986).
@@
By Taylor’s theorem with integral remainder (Folland,, 2005), for each subset index $\mathcal{I}_{j}$,
@@
The Lipschitz conditions are to ensure that the SDE has a unique strong solution with uniformly bounded moments. It is possible to appropriately relax them if we allow weak solutions (Mil’shtein,, 1986).
@@ Appendix D Auxiliary Lemmas
Theorem D.1 (Existence and Uniqueness of Strong Solutions (Evans,, 2012)).
@@
(Li et al.,, 2019) Let $\{x_{k}:k\geq 0\}$ be the generalized iterations defined in (22). Suppose
@@
We will use the following lemma on convex order to prove Proposition 3.9, whose proof can be found in classical textbooks on stochastic order, such as Müller and Stoyan, (2002).
@@ Appendix F Additional Related Works
Theoretical understanding of SAM. Although SAM and its variants have achieved remarkable success in various practical applications (Foret et al.,, 2021; Kwon et al.,, 2021; Kaddour et al.,, 2022; Li et al., 2024b, ; Li and Giannakis,, 2024), the theoretical understanding behind them remains limited. The pioneering work of Andriushchenko and Flammarion, (2022) provided the first theoretical framework for understanding SAM, covering its convergence properties and implicit bias in simple network structures, while also systematically illustrating several empirical phenomena. Subsequently, Si and Yun, (2024) extended the convergence analysis to various deterministic and stochastic settings. Compagnoni et al., (2023); Luo et al., (2025) conducted an in-depth analysis of the dynamics of SAM using the SDE framework previously developed by Li et al., (2017), leading to a deeper understanding of its implicit bias. On the other hand, Wen et al., (2022) investigated the implicit bias of SAM by analyzing its slow ordinary differential equation (ODE) behavior near the minimizer manifold, demonstrating how SAM drifts toward flatter minima. More recently, Zhou et al., (2024) studied the late-stage behavior of SAM using stability analysis, showing its advantage in escaping sharp minima.
@@ Appendix I Algorithm: Reweighted SAM
We present the experimental results of applying the proposed reweighting strategy to ASAM (Kwon et al.,, 2021) in Table 9. The results demonstrate consistent improvements, and further investigation into the effectiveness of applying the reweighting strategy to different SAM variants remains an interesting direction for future work.

### v3

@@ 5.1 Training from scratch
We evaluate three optimization methods SGD, SAM [^2] In this section, consistent with common practice, we use “SAM” to refer to the mini-batch SAM., and RW-SAM on CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009), training three models from scratch: ResNet-18, ResNet-50 (He et al., 2016), and WideResNet-28-10 (Zagoruyko and Komodakis, 2016). We use a batch size of 128 and a cosine learning rate schedule with an initial learning rate of 0.1. SAM and RW-SAM are trained for 200 epochs, while SGD is trained for 400 epochs. We apply a momentum of 0.9 and a weight decay of $5e^{-4}$, along with standard data augmentation techniques, including horizontal flipping, padding by four pixels, and random cropping.
@@
For large-scale experiments, we train a ResNet-50 on ImageNet-1K (Deng et al., 2009) for 90 epochs with an initial learning rate of 0.05. For both SAM and RW-SAM, we use the same $\rho=0.05$. For the additional hyperparameter $\lambda$ in RW-SAM, we set $\lambda=0.25$. All other hyperparameters remain the same as those used on CIFAR-10/100.
@@
Analysis of Computational Overhead. RW-SAM requires an additional forward pass to estimate per-sample gradient norms, leading to approximately 1/6 more training overhead compared to vanilla SAM, as a forward pass typically takes about half the time of a backward pass (Kaplan, 2022). For a report of the additional wall‐clock time overhead, please see Table 8 in Appendix J. However, RW-SAM matches the performance of m-SAM at $m=64$ without incurring its nearly two-fold training overhead (See Table 7 in Appendix J). This highlights the efficiency of RW-SAM in balancing computational cost and performance.
@@ 5.2 Fine-tuning
We fine-tune a ViT-B/16 model (Dosovitskiy et al., 2020), pre-trained on ImageNet-1K, on CIFAR-10 and CIFAR-100. We train for 20 epochs with an initial learning rate of 0.01. Other hyperparameters are the same as those in training from scratch. The results are summarized in Table 2(b). We also fine-tune a pretrained DistilBERT model (Sanh et al., 2019) on the GLUE benchmark (Wang et al., 2018). We use AdamW (Loshchilov and Hutter, 2019) as the base optimizer. The detailed hyperparameter settings are provided in Table 10 of Appendix J, and the results are presented in Table 3. We observe that RW-SAM consistently outperforms SAM in both fine-tuning experiments.
@@ 5.3 Robustness to label noise
Foret et al. (2021) have shown that SAM exhibits robustness to label noise. Motivated by this, we evaluate RW-SAM’s performance on CIFAR-10 with labels randomly flipped at specified noise ratios. We train a ResNet-18 and report clean test accuracies in Table 4. As the noise ratio increases, RW-SAM maintains remarkably strong performance. In particular, at an 80 % noise ratio, RW-SAM achieves a 16 % absolute accuracy improvement over SAM.
@@ Acknowledgement
Following the SDE framework by Mil’shtein (1986); Li et al. (2017); Compagnoni et al. (2023); Luo et al. (2025), we have the following definition:
@@
(Adaption of Theorem 3 in Li et al. (2019)) Let $T>0$, $\eta\in(0,\min\{1,T\})$, $N=\lfloor T/\eta\rfloor$. Let $\alpha\geq 1$ be an integer. Suppose further that the following conditions hold:
@@
By substituting Assumption (i) in the penultimate step of the proof in Theorem 3 in Li et al. (2019) with Assumption (i) of our Theorem A.2, the same argument goes through and yields the desired conclusion. Hence, we omit the proof.
Next, we state the key lemma for the two-parameter weak approximation. By employing a Dynkin expansion (see, e.g. Evans (2012)) instead of a full Itô–Taylor expansion, it avoids the proliferation of terms and provides a streamlined approach to controlling the remainder error in the two-parameter setting.
@@ B.1 Mini-batch USAM
Theorem B.2 (mini-batch USAM SDE, adapted from Theorem 3.2 of Compagnoni et al. (2023)).
@@
Theorem B.2 follows by replacing the single‐parameter Lemmas A.1, A.2 and A.5 in Compagnoni et al. (2023) with our two‐parameter versions—Theorem A.2, Lemma A.4 and Lemma A.5—imposing the extra global Lipschitz conditions to guarantee existence and uniqueness of the strong solution, and using our Assumption 3.1 to expand the drift term. We therefore omit the routine algebraic details. ∎
@@
By Taylor’s theorem with integral remainder (Folland, 2005) we have, for each $\gamma$,
@@
The Lipschitz conditions are to ensure that the SDE has a unique strong solution with uniformly bounded moments. It is possible to appropriately relax them if we allow weak solutions (Mil’shtein, 1986).
@@
By Taylor’s theorem with integral remainder (Folland, 2005) we have, for each $\gamma$,
@@
The Lipschitz conditions are to ensure that the SDE has a unique strong solution with uniformly bounded moments. It is possible to appropriately relax them if we allow weak solutions (Mil’shtein, 1986).
@@ C.1 Mini-batch SAM
Theorem C.2 (mini-batch SAM SDE, adapted from Theorem 3.5 of Compagnoni et al. (2023)).
@@
    $\nabla f^{\mathrm{SAM}}\;=\;\nabla\bigl(f+\rho\mathbb{E}\|\nabla f_{\gamma}\|\bigr)\quad\text{and}\quad\sqrt{\Sigma^{\mathrm{SAM}}}$
@@
Theorem C.2 follows by replacing the single‐parameter Lemmas A.1, A.2 and A.14 in Compagnoni et al. (2023) with our two‐parameter versions—Theorem A.2, Lemma A.4 and Lemma A.5—imposing the extra global Lipschitz conditions to guarantee existence and uniqueness of the strong solution. We therefore omit the routine algebraic details. ∎
@@
By Taylor’s theorem with integral remainder (Folland, 2005), for each $\gamma$ and writing $u(x)=\nabla f(x)/\|\nabla f(x)\|$, we have
@@
The Lipschitz conditions are to ensure that the SDE has a unique strong solution with uniformly bounded moments. It is possible to appropriately relax them if we allow weak solutions (Mil’shtein, 1986).
@@
By Taylor’s theorem with integral remainder (Folland, 2005), for each subset index $\mathcal{I}_{j}$,
@@
The Lipschitz conditions are to ensure that the SDE has a unique strong solution with uniformly bounded moments. It is possible to appropriately relax them if we allow weak solutions (Mil’shtein, 1986).
@@ Appendix D Auxiliary Lemmas
Theorem D.1 (Existence and Uniqueness of Strong Solutions (Evans, 2012)).
@@
(Li et al., 2019) Let $\{x_{k}:k\geq 0\}$ be the generalized iterations defined in (22). Suppose
@@
We will use the following lemma on convex order to prove Proposition 3.9, whose proof can be found in classical textbooks on stochastic order, such as Müller and Stoyan (2002).
@@ Appendix F Additional Related Works
Theoretical understanding of SAM. Although SAM and its variants have achieved remarkable success in various practical applications (Foret et al., 2021; Kwon et al., 2021; Kaddour et al., 2022; Li et al., 2024b; Li and Giannakis, 2024), the theoretical understanding behind them remains limited. The pioneering work of Andriushchenko and Flammarion (2022) provided the first theoretical framework for understanding SAM, covering its convergence properties and implicit bias in simple network structures, while also systematically illustrating several empirical phenomena. Subsequently, Si and Yun (2024) extended the convergence analysis to various deterministic and stochastic settings. Compagnoni et al. (2023); Luo et al. (2025) conducted an in-depth analysis of the dynamics of SAM using the SDE framework previously developed by Li et al. (2017), leading to a deeper understanding of its implicit bias. On the other hand, Wen et al. (2022) investigated the implicit bias of SAM by analyzing its slow ordinary differential equation (ODE) behavior near the minimizer manifold, demonstrating how SAM drifts toward flatter minima. More recently, Zhou et al. (2024) studied the late-stage behavior of SAM using stability analysis, showing its advantage in escaping sharp minima.
@@ Appendix I Algorithm: Reweighted SAM
We present the experimental results of applying the proposed reweighting strategy to ASAM (Kwon et al., 2021) in Table 9. The results demonstrate consistent improvements, and further investigation into the effectiveness of applying the reweighting strategy to different SAM variants remains an interesting direction for future work.